OpenStudy (anonymous):

What's the inverse of: ln(2x)-3. My book says it's 1/2e^(x+3)

OpenStudy (anonymous):

the guy showed you in your last post how to do it.. did you understand it?

OpenStudy (anonymous):

it gives you $$\frac{e^{x+3}}{2}$$ which is equal to $$\frac{1}{2}e^{x+3}$$

OpenStudy (anonymous):

How come the +3 is in the exponent

OpenStudy (anonymous):

what is ln(2x)-3 =? is it y

OpenStudy (anonymous):

you can assign it to any variable, conventionally that variable would be y

OpenStudy (anonymous):

may be it is ln(2x)-3=x then u get the ur answer @minisweet4

OpenStudy (phi):

you start with y = ln(2x)-3 rename Y to X and vice versa: x = ln(2y)-3 add 3 to both sides x+3 = ln(2y) or ln(2y)= (x+3) now the "hard part". to "undo the ln" make both sides the exponent of the base "e"

OpenStudy (anonymous):

yes this is exactly what the last guy showed, except he swapped the x and y at the end

OpenStudy (anonymous):

this can be done due to the properties of logs and because lnx = log base e of x

OpenStudy (anonymous):

to undo the ln